OpenStudy (anonymous):

The rate traveled from Amarillo to Austin by a bus averages 60 miles per hour. The bus arrived in Austin after eight hours of travel. An automobile averages 48 miles per hour. Using the inverse variation relationship, show what the time would be for the automobile to complete the trip. i got 12 hours? is that right?

OpenStudy (anonymous):

60*8=480 480/48=10

OpenStudy (saifoo.khan):

\[Inverse-variation-in- algebra.?\]

OpenStudy (saifoo.khan):

I believe this is y=k/x where k is your constant. x = 60 y = 8 8 = k/60 k = 480 Now that you have the value for k, you can solve the second part. y = k/x k = 480 x = 48 y = 480/48 y = 10 hours.

OpenStudy (anonymous):

sorry to intrude but can someone help me with finance on here?

OpenStudy (gw2011):

(60)(8)=480 miles traveled You can use the distance formula: d=(r)(t) where d=distance where r=rate where t=time 480=(48)(t) 480/48=t t=10 hours